OpenStudy (anonymous):

How do i plot r^2=-sin(theta)

OpenStudy (anonymous):

find table of values for r with corresponding value of theta......... plot r in y axis and theta in x axis

OpenStudy (anonymous):

but if i take the square root of both sides, i can have that negative sin under the square root can i

OpenStudy (anonymous):

ya you cant have negative sign under root........for the values of theta greater than 180 and less than 360 , you will have negative values of sintheta so negative and negative makes positive..........

OpenStudy (anonymous):

What about, how do i test for symmetry

OpenStudy (anonymous):

what does test for symmetry mean exactly, i am not getting it....... you told previously that you need to plot on the graph

OpenStudy (anonymous):

No can you help me with this, i am just learning how to do polar equations if you could hsow me how to do this one it would greatly help

OpenStudy (anonymous):

Well i know r=a is a line, would r^2 be a parabola

OpenStudy (anonymous):

2bornoto2b, you are quite a teacher

OpenStudy (anonymous):

yes ma'am

OpenStudy (anonymous):

yeah, i did thank you, But in my book the graph is of two circles

OpenStudy (anonymous):

2bornoto2b, what software you use?

OpenStudy (anonymous):



OpenStudy (anonymous):

2bornto2b, why we need to slipt the graph............

OpenStudy (anonymous):

Something like that, but the two circle are symmetric

OpenStudy (2bornot2b):

Oh dear, then its a polar equation. Its not a coordinate one.

OpenStudy (anonymous):

yeah, its a polar equation, sorry i forget to mention

OpenStudy (anonymous):

oh.......but he has not mentioned that

OpenStudy (2bornot2b):

Actually I guessed it because he used to variables, and I changed them to one, silently

OpenStudy (2bornot2b):

*two not to

OpenStudy (anonymous):

oh! but what about the diagram he has shown

OpenStudy (2bornot2b):

I don't have any idea about your level of education. But still I guess you know these x=rcos@, y=rsin@ r=(x^2+y^2)^0.5 Do you know these?

OpenStudy (anonymous):

yes

OpenStudy (2bornot2b):

Now do you realize that the equation in coordinate form is (x^2+y^2)^(3/2)=-y

OpenStudy (anonymous):

I take that back

OpenStudy (anonymous):

What? how i that possible

OpenStudy (anonymous):

is*

OpenStudy (anonymous):

I understand thje x^2+y^2 but how come it is ^3/2

OpenStudy (2bornot2b):

Use pen and paper, and use the relations x=rcos@, y=rsin@ r=(x^2+y^2)^0.5

OpenStudy (anonymous):

yea, i get that one, but what about the ^(3/2) where is that coming from

OpenStudy (2bornot2b):

1/2+1

OpenStudy (anonymous):

why are you adding a one

OpenStudy (2bornot2b):